OpenStudy (anonymous):

Question about an absolute value equation I got wrong and I don't know why. 3 I 2a-5 I = 21

OpenStudy (anonymous):

|2a - 5| = 7 2a - 5 = 7 ; 2a - 5 = -7 2a = 12; 2a = -2 a = 6, -1

OpenStudy (anonymous):

That is exactly what I got but the A+dvancer math program I am using said it was wrong. Maybe I did a typo...?

OpenStudy (anonymous):

I am absolutely positive those answers are correct. You can plug them in to check. They are also the only two possible answers if you construct the graph.